OpenStudy (anonymous):

Determine the number & type of solutions for x^2+4x+3=0 a) 1 rational double b) 2 irrational c) 2 rational d) 2 complex Someone please help me

OpenStudy (anonymous):

Try solving it using the quadratic formula -b +or- squareroot( b^2 - 4ac) / 2a

OpenStudy (anonymous):

where a =1 b =4 and c=3 in this case.

OpenStudy (anonymous):

Or you can simplify x^2+4x+3 x^2+4x+3= (x + 1)(x + 3)

OpenStudy (anonymous):

in this case you can see that x will equal -3 and -1.

OpenStudy (anonymous):

I basically solved it for you :/ all you have to do is figure out if -3 and -1 are rational, irrational or complex

OpenStudy (anonymous):

for a quadratic equation, as Romero says, we have a formula the important bit for evaluating the type of root is called the "discriminant" its the bit in the square root if \[ax^2 + bx + c =0\] the discriminant is \[\Delta = b^2 - 4ac\] if \[\Delta = n^2\] where n is an integer we get rational roots if \[\Delta < 0\] we have complex roots if \[\Delta = 0\] we have repeated roots etc etc

OpenStudy (anonymous):

they are rational, right?

OpenStudy (anonymous):

Yes, they are rational numbers.

OpenStudy (anonymous):

thank you everyone.